# Simultaneous symmetry breaking in spontaneous Floquet states: Floquet-Nambu-Goldstone modes, Floquet thermodynamics, and the time operator

### v2

@@ A.2 GPG solutions in a ring
Since the GPG equation can be rewritten as Eq. (84), we first study in detail the solutions to the stationary GP equation
@@
These densities are related to the more physical chemical potential $\mu$, the particle current $J$, and the amplitude energy $E_{A}$ as
@@
Moreover, $|\Psi_{0}(x)|^{2}=|\Psi_{\mathbf{n}}(x)|^{2}$, and thus the total particle number is directly obtained from the integral of the cnoidal density. With the help of these results, one arrives at Eq. (93). The trick to solve that equation is to express all roots $n_{i}$ in terms of $\nu$, $n_{i}=n_{i}(\nu)$, achieving thus an equation for the real parameter $\nu$. For that purpose, one combines the first two lines of Eq. (93) to express $n_{1},n_{3}$ in terms of $\nu$, and invokes the definition of $\nu$ itself, Eq. (86), to obtain $n_{2}(\nu)$. Inserting these results into the third line of Eq. (93) yields a single equation for $\nu$, which can be easily numerically solved as $0\leq\nu\leq 1$. Once $\nu$ is obtained, the densities $n_{i}=n_{i}(\nu)$ are computed, and the complete GPG wavefunction and the total energy are determined.
Finally, we show that the Berry-Gibbs connection (53) automatically vanishes in our gauge choice. By decomposing the GPG wavefunction in phase and amplitude, $\Psi_{0}(\mathbf{x})=\sqrt{n(\mathbf{x})}e^{i\eta(\mathbf{x})}$, it is easily seen that in general

### v3

@@ A.2 GPG solutions in a ring
Since the GPG equation can be rewritten as Eq. (84), we first study in detail the solutions to the stationary homogeneous GP equation
@@
These densities are related to the more physical chemical potential $\mu$, particle current $J$, and amplitude energy $E_{A}$ as
@@
Moreover, $|\Psi_{0}(x)|^{2}=|\Psi_{\mathbf{n}}(x)|^{2}$, and thus the total particle number is directly obtained from the integral of the cnoidal density. With the help of these results, one arrives at Eq. (93). The trick to solve that system of equations is to express all roots $n_{i}$ in terms of $\nu$, $n_{i}=n_{i}(\nu)$. For that purpose, one combines the first two lines of Eq. (93) to express $n_{1},n_{3}$ in terms of $\nu$, and invokes the definition of $\nu$ itself, Eq. (86), to obtain $n_{2}(\nu)$. Inserting these results into the third line of Eq. (93) yields a single equation for $\nu$, which can be easily numerically solved as $0\leq\nu\leq 1$. Once $\nu$ is obtained, the densities $n_{i}=n_{i}(\nu)$ are computed, and the complete GPG wavefunction and the total energy are determined.
Finally, we show that the Berry-Gibbs curvature (53) automatically vanishes in our gauge choice. By decomposing the GPG wavefunction in phase and amplitude, $\Psi_{0}(\mathbf{x})=\sqrt{n(\mathbf{x})}e^{i\eta(\mathbf{x})}$, it is easily seen that in general